Grade 12Electrostatics

an electron of mass M,initially at rest,moves through a certain distance in a uniform electric field in time T.A proton of mass m,also initially at rest, takes time t to move through an equal distance in this uniform electric field.Negelecting the effect of gravity,the ratio of t/T is nearly equal to...??

Ashik Rahman
12 Years agoGrade 12

2 Answers

Arun Kumar
12 Years ago

Acceleration inversely proportional to mass.
Thanks & Regards
Arun Kumar
IIT Delhi
Askiitians Faculty
Kushagra Madhukar
6 Years ago
Dear student,
Please find the attached solution to your problem below.
 
Let the subscript of the term represent the particle.
Now, distance travelled by proton and electron is same under given constraints. (xp = xe)
Hence,
Hence, t/T = te/tp = (M/m)½ 
 
Hope it helps.
Thanks and regards,
Kushagra